Please help me solve:
(2t-5)^2=24
I know I have to use the quadratic formula, but I can't figure it out.

Answer by stanbon(75887)

Please help me solve:
(2t-5)^2=24
---
4t^2 - 20t + 25 = 24
---
4t^2 - 20t + 1 = 0
t = [20 +- sqrt(400 - 4*4*1)/(8)
---
t = [20 +- sqrt(384)]/8
---
t = [20 +- 19.6]/8
---
t = 39.6/8 or t = 0.4/8
---------------------------
Cheers,
Stan H.